OpenStudy (anonymous):

Problem Set 1: IA 4 b and c: I'm confused on even and odd equations, I guess. Can anyone further explain what the difference is? (... They're dividing by two to get f(x) instead of 2f(x) in b, I get that, mostly, but in c, how are they finding out if they're even or odd? Is it just intuitive b/c x cubed of an odd will = an odd number?)

OpenStudy (anonymous):

it's very simple just replace every (x) with (-x) if it's the same equation (f(-x)=f(x)) then it's even if it equals negative the original equation (f(-x)=-f(x)) then it's odd

OpenStudy (phi):

x cubed of an odd will = an odd number? that motivates the names: contrast the graph of x^2 and x^3 notice for y= x^2 for x= 1 or -1 we get y=1 or , in general, for \( x= \pm x_0\) , y = \( x_0^2\) we can write this as f(x) = f(-x) and this is the definition of an even function. similarly, f(x) = - f(-x) is the definition of an odd function.

OpenStudy (anonymous):

Functions are classified into even/odd for a better understanding of their symmetry. If a function of x is even, then its graph on the Cartesian co-ordinates would be symmetric about the Y axis; i.e. the Y axis splits the curve into two equal and congruent halves. For example if f(2) and f(-2) exist then this would imply, f(2) = f(-2). For any general value of x, check if f(x) = f(-x) OR f(x) - f(-x) = 0 If a function of x is odd then its graph would be symmetric about the origin; i.e. there will be equal and congruent halves of the curve in opposite quadrants. For example if f(1) and f(-1) exist then according to the symmetry condition it can be said that f(-1) = - f(1). For any general value of x, check if f(-x) = -f(x) OR f(x)- f(-x) = 0